QUESTION 19 Which one of the following direct jumps in hydrogen atom results in emission of a photon (O A jump from sixth level to the second level A jump from the third level to the second level O A jump from the fourth level to the second level a jump from the fifth level to the second level QUESTION 20 An object has 14.3x1019 more electrons. What is the charge of this object in coulombs QUES TION 21 What is the current through a 444? resistor in amperes, if a 299V voltage is applied a
Answer #1

20.

Q = n*e

n = excess electrons

e = charge on electron

Q = 14.3*10^19*(-1.6*10^-19)

Q = -22.88 C

21

Using ohm's law:

V = i*R

i = V/R = 29.9/44.4

i = 0.673 Amp

Please Upvote.